OpenStudy (richyw):

Help with Green's theorem/vector calculus identities My notes say I can get Green's Theorem \[\int_V \left(w \nabla^2 u - u \nabla^2w\right)dV = \int_S \left( w\nabla u - u\nabla w\right)\cdot d\mathbf{S}\] by substituting \(\mathbf{A}=w\nabla u\) into the divergence theorem \[\int_V \left(\nabla\cdot\mathbf{A}\right)dV = \int_S \mathbf{A}\cdot d\mathbf{S}\]

OpenStudy (richyw):

I am assuming that I am missing some vector calculus identity here.

OpenStudy (irishboy123):

if you are writing the div theorem as \(\int_V \nabla \bullet \vec A ~ dV = \int_S \vec A \bullet d \vec S\) and set \(\vec A = w \nabla u\) then sub that into div theorem \(\int_V \nabla \bullet (w \nabla u) ~ dV = \int_S (w \nabla u) \bullet d \vec S\) \(\int_V \nabla w \bullet \nabla u + w \nabla^2 u ~ dV = \int_S (w \nabla u) \bullet d \vec S \qquad \star \) then switch u and w around to get \(\int_V \nabla u \bullet \nabla w + u \nabla^2 w ~ dV = \int_S (u \nabla w) \bullet d \vec S \qquad \circ\) then \(\star - \circ \implies\) \( \int_V \nabla w \bullet \nabla u + w \nabla^2 u - (\nabla u \bullet \nabla w + u \nabla^2 w) ~ dV = \int_S (w \nabla u - u \nabla w )\bullet d \vec S \qquad \) \( \int_V w \nabla^2 u - u \nabla^2 w ~ dV = \int_S (w \nabla u - u \nabla v )\bullet d \vec S \qquad \)